OpenStudy (anonymous):

The pair of points are inversely related. Find the missing variable. (9, a) and (6, 12)

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

inverse variation \(\large\tt \begin{align} \color{black}{y=\dfrac{k}{x}\\~\\ a=\dfrac{k}{9}\\~\\ 12=\dfrac{k}{6}\\~\\ }\end{align}\) first find the vaue of k and then a